A source produces first-order lines when incident normally on a 9000 slits/cm diffraction grating at angles θ1 = 28.8°, θ1 = 36.7°, θ1 = 38.6°, and θ1 = 41.2° 


Part A 

What are the wavelengths? 

Express your answers using three significant figures separated by commas.  
Answer #1

m = 1 order

Slit separation = D

Dsin\theta = m\lambda

D = 1/N

N = lines/cm = 9000

\lambda = Dsin\theta/m

\lambda = (sin\theta)/(N*m)

\lambda1 = (sin28.8)/(9000*1)

\lambda1 = 5.3528 x 10-5 cm

\lambda1 = 535 nm

\lambda2 = (sin36.7)/(9000*1)

\lambda2 = 6.6402 x 10-5 cm

\lambda2 = 664 nm

\lambda3 = (sin38.6)/(9000*1)

\lambda3 = 6.9319 x 10-5 cm

\lambda3 = 693 nm

\lambda4 = (sin41.2)/(9000*1)

\lambda4 = 7.3187 x 10-5 cm

\lambda4 = 731 nm